OpenStudy (anonymous):

Mary's paycheck was for $600.00. When she cashed her check at the bank, the teller gave her a total of 40 bills, both tens and twenties. How many tens did the teller give Mary?

OpenStudy (anonymous):

Think you can help me??

OpenStudy (shubhamsrg):

let no. of 10's be x and no. of 20's be y now given that x+y = 40 also 10x + 20y = 600 just solve for x & y.

OpenStudy (anonymous):

huh??

OpenStudy (shubhamsrg):

x+y =40 and 10x + 20y = 600 solve them simultaneously.. x is number of tens//